Another method for graphing a line is by finding the x and y-intercepts. Explain how you find these intercepts.
1
Expert's answer
2020-02-26T10:22:53-0500

x-intercept and y-intercept of a line are the points where the line intersect/crosses the x-axis and y-axis respectively.



We also know on x-axis, values of y coordinates are zero (in another words, y=0 represent x-axis). Similarly on y-axis, values of x coordinates are zero (in another words , x=0 represent y-axis).


So for finding x-intercept, we substitute y=0 in the given equation of line. Similarly for finding y-intercept, we substitute x=0 in the given equation of line.


I will also explain this with an example. For example consider we want to graph the line x + 2y = 4

For finding x -intercept, we put y=0 in above equation

x + 2(0) = 4

x + 0 = 4

x = 4

Thus (4, 0) is the x- intercept.

Similarly for finding y-intercept, we put x = 0 in the equation of line.

0 + 2y = 4

2y = 4

y = 2

Thus (0, 2) is the y-intercept.

For graphing equation of line we mark these two coordinates on x-axis and y-axis respectively and draw a line passing through them as shown below.